### v1

# Holmes: Towards Distributed Training Across Clusters with Heterogeneous NIC Environment

Abstract
Large language models (LLMs) such as GPT-3, OPT, and LLaMA have demonstrated remarkable accuracy in a wide range of tasks. However, training these models can incur significant expenses, often requiring tens of thousands of GPUs for months of continuous operation. Typically, this training is carried out in specialized GPU clusters equipped with homogeneous high-speed Remote Direct Memory Access (RDMA) network interface cards (NICs). The acquisition and maintenance of such dedicated clusters is challenging. Current LLM training frameworks, like Megatron-LM and Megatron-DeepSpeed, focus primarily on optimizing training within homogeneous cluster settings. In this paper, we introduce Holmes, a training framework for LLMs that employs thoughtfully crafted data and model parallelism strategies over the heterogeneous NIC environment. Our primary technical contribution lies in a novel scheduling method that intelligently allocates distinct computational tasklets in LLM training to specific groups of GPU devices based on the characteristics of their connected NICs. Furthermore, our proposed framework, utilizing pipeline parallel techniques, demonstrates scalability to multiple GPU clusters, even in scenarios without high-speed interconnects between nodes in distinct clusters. We conducted comprehensive experiments that involved various scenarios in the heterogeneous NIC environment. In most cases, our framework achieves performance levels close to those achievable with homogeneous RDMA-capable networks (InfiniBand or RoCE), significantly exceeding training efficiency within the pure Ethernet environment. Additionally, we verified that our framework outperforms other mainstream LLM frameworks under heterogeneous NIC environment in terms of training efficiency and can be seamlessly integrated with them.
1 Introduction
In recent years, the emergence of large language models (LLMs) [33, 42] such as GPT-3 [22], OPT [41], and LLaMA [34] has significantly transformed the landscape of natural language processing (NLP) by achieving state-of-the-art accuracy in a wide variety of tasks. Despite these remarkable advancements, LLM training is often accompanied by considerable costs, requiring the sustained operation of tens of thousands of advanced GPUs over extended periods. For instance, Meta’s 175 billion parameter OPT-175 was trained for 33 days on 1024 NVIDIA A100 GPUs. A slightly smaller model, LLaMa with 65 billion parameters, was trained for 21 days using a cluster of 2048 NVIDIA A100 GPUs [41, 34, 22, 40]. Conventionally, these models undergo training within specialized clusters that are equipped with homogeneous high-speed Remote Direct Memory Access (RDMA) network interface cards (NICs). However, the acquisition and maintenance of such dedicated clusters can present substantial financial and logistical challenges [34, 42].
Figure 1: An overview of the Holmes framework. Holmes provides training support for a diverse range of LLM types and seamlessly integrates with contemporary mainstream LLM training frameworks. This flexibility extends to the training of LLMs in the heterogeneous NIC environment, where Holmes consistently achieves performance levels that are close to those achievable in the homogeneous NIC settings.
RDMA technology has found extensive deployment in modern data centers [20, 11], offering low latency and high throughput benefits that are particularly advantageous for LLM training. RDMA technology primarily comprises two categories: InfiniBand and RoCE (RDMA over Converged Ethernet) [7, 13, 11]. Both of these technologies facilitate direct access to remote computer memory, bypassing the need for CPU intervention, thus leading to more efficient data transfer and communication. InfiniBand represents a dedicated networking technology, while RoCE is a protocol designed to implement RDMA within the Ethernet framework. Both InfiniBand and RoCE play an integral role in the construction of high-performance computing and data center networks [17, 13, 11]. Despite their widespread applications, the two networking technologies, InfiniBand and RoCE, are inherently incompatible with each other. This incompatibility poses challenges, especially in the context of training LLMs within clusters that incorporate various RDMA NICs. In such scenarios, the adoption of pure Ethernet for communication often emerges as the only viable option.
Speeding up the LLM training process and minimizing associated costs have been persistent areas of research. Current LLM training frameworks, as evidenced in recent surveys [42], which feature prominent models like Megatron-LM [32, 25], DeepSpeed [28, 33, 1], and FairScale [9], predominantly prioritize speed within homogeneous data center networks. Consequently, these frameworks may not be easily adaptable to the diverse landscape of the heterogeneous RDMA NIC environment. To address these challenges, there is a need for an open and unified LLM training framework that can seamlessly accommodate the heterogeneity of RDMA NICs across various clusters, thus alleviating the operational complexity associated with large-scale datacenter networks.
In this paper, we present Holmes, a high-performance framework optimized for training LLMs in the heterogeneous RDMA NIC environment. An overview of the framework is illustrated in Figure 1. Holmes optimally leverages the diverse capabilities of NICs within cluster configurations. A key contribution lies in our intelligent scheduling methodology, strategically allocating computational tasklets to GPUs based on their associated NIC types. This approach aligns computational tasks with the strengths of each network technology, including InfiniBand, RoCE, and Ethernet. As a result, Holmes significantly mitigates overall training costs.
Specifically, our framework tracks hardware topology and network characteristics, dynamically assigning parallel work groups to suitable GPU devices during LLM training. The internal scheduling method, which is applicable to the heterogeneous NIC environment, is complemented by well-designed optimizations in both data and model parallelism [38, 15, 23, 21]. As a result, Holmes can effectively harness the full potential of the real-world production infrastructure. This includes the use of high-speed RDMA NICs for data parallelism, which exhibits a high communication overhead [21, 27, 30], and Ethernet NICs for pipeline model parallelism, known for its lower communication overhead [15]. We demonstrate impressive performance in using our framework to train LLM across clusters equipped with heterogeneous NICs.
The proposed Holmes framework demonstrates exceptional scalability through pipeline model parallelism, seamlessly extending to multiple GPU clusters. This scalability is maintained even in scenarios where high-speed interconnects (InfiniBand or RoCE) are unavailable between nodes from distinct clusters. Therefore, Holmes can be applied to train LLMs across multiple GPU clusters at different locations. This maximizes the utilization of existing GPU clusters for LLM training, eliminating the need to reconstruct larger GPU clusters to accommodate numerous GPU devices and RDMA NICs specifically for training purposes. To validate Holmes’s scalability, we conducted an extensive range of experiments covering a diverse spectrum of scenarios involving heterogeneous NICs. Remarkably, our training framework consistently achieved performance levels close to those attainable in environments with homogeneous RDMA-capable networks, such as pure InfiniBand or RoCE. Furthermore, Holmes notably surpassed training efficiency even in the pure Ethernet environment.
In general, Holmes provides a highly optimized framework for training LLMs in the heterogeneous RDMA NIC environment. Our optimizations of schedule and parallelism unlock the potential of heterogeneous systems for efficient and scalable LLM training. We firmly believe that Holmes signifies a significant step in democratizing LLMs, making efficient scaling attainable for a broader research community.
Our contributions are summarized as follows:
  • We propose Holmes, a novel LLM training framework designed to adapt to the heterogeneous NIC environment. Holmes is non-intrusive to existing infrastructures, harnessing pipeline model parallelism technologies. It facilitates LLM training across multiple GPU clusters without the need to reconstruct larger GPU clusters. This feature is crucial for conserving computational resources.
  • Moreover, Holmes introduces a method for tailoring stage division in pipeline model parallelism. This method takes into account the speed variations inherent in heterogeneous NICs and allocates varying numbers of layers to different stages accordingly.
  • We conducted comprehensive experiments to demonstrate that training LLMs with Holmes can achieve performance close to those achievable with homogeneous RDMA-capable networks, significantly exceeding training efficiency within the pure Ethernet environment.
  • Furthermore, Holmes seamlessly integrates with other mainstream LLM training frameworks such as Megatron-LM and Megatron-DeepSpeed.
Limitations and Future Directions. In this paper, we focus on a fundamental part of LLM training, but there are important practical aspects that we have not yet covered. We assume that communication between devices is stable and that all devices are consistently online. It is important to note that we have not fully trained the system to convergence; instead, we have performed partial training to validate our approach. In the future, we need to explore scheduling methods for diverse environments and figure out how to handle faults.
2 Problem Formulation
In this section, we first provide an overview of the background and motivations behind our work. Following this, we introduce different NIC environments. Finally, we describe the problem addressed in this paper and proceed to formalize it.
2.1 Background and Motivation
2.1.1 RDMA Preliminaries
RDMA is a technology that enables super-efficient data exchange between computers on a network. It allows one computer to access the memory of another without bothering the CPU of either machine. This direct memory access results in lightning-fast data transfers and reduced delay. RDMA substantially improves network performance and efficiency, making it indispensable for applications that demand high-speed data transmission and rapid, low-latency communication. It is particularly crucial in domains such as high-performance computing, cloud computing, and expansive data centers. Two common ways to implement RDMA are InfiniBand and RoCE. InfiniBand is a specialized networking technology, whereas RoCE adapts RDMA to work over Ethernet networks. Both approaches facilitate direct memory access between computers, streamlining data transfer, and communication. The choice between them depends on factors such as network design, specific needs, and the context of the application.
The impact of RDMA technology on LLM training is significant, manifesting in various aspects:
  • Faster Communication Speed: RDMA’s direct memory access accelerates communication operations like model parameter updates and gradient aggregation in distributed environments.
  • More Efficient Data Transfer: RDMA facilitates efficient data transfer, crucial for transmitting parameters and gradients in LLMs. In scenarios with substantial data, efficient transfer enhances training performance.
  • Better Horizontal Scalability: RDMA enhances horizontal scalability, enabling effective LLM training in large GPU clusters without undue concerns about communication bottlenecks.
2.1.2 LLM Infrastructure
Training LLMs requires an advanced infrastructure, including GPU clusters, high-speed RDMA NICs, and ample storage capacity. In particular, the demand for NVIDIA A100 GPU devices often exceeds 5,000 [25]. As LLM continues to grow in size, this demand is expected to increase further. However, the development of cluster infrastructure has lagged behind rapid advances in LLM. Many existing GPU clusters, constructed several years ago, lack the capacity to accommodate more than 1,000 NVIDIA A100 GPU devices, and network connectivity between GPU clusters in different locations is often limited. Consequently, the need to reconstruct larger GPU clusters for LLM research has become a necessity.
Moreover, the RDMA technology has two implementations: InfiniBand and RoCE. Different GPU clusters may be equipped with different types of RDMA NIC, and these two variants are not mutually compatible. Existing training frameworks, such as Megatron-LM [32, 25] and Megatron-DeepSpeed [28, 1], face challenges in training LLM in the heterogeneous NIC environment. Consequently, they are unable to fully harness GPU devices in the heterogeneous clusters. This limitation leads to a suboptimal utilization of GPU computing resources.
2.1.3 Motivation
These practical challenges in infrastructure prompt us to reconsider LLM training, focusing on the effective utilization of GPU devices and RDMA NICs across multiple clusters. Our objective is to develop a unified LLM training framework with the following key goals:
  • Compatibility: The training framework should seamlessly operate with different RDMA NICs across multiple GPU clusters. To accomplish this, we redesign existing schemes for parallelism methods.
  • Scalability: The training framework should easily scale to accommodate thousands of GPU devices. For this purpose, we have developed Holmes based on the Megatron-LM framework, a popular LLM training framework known for successfully handling training iterations on models with up to a trillion parameters.
  • Availability: Modern large-scale GPU clusters are consistently equipped with RDMA NICs. Our training framework maximizes the use of these available machine resources while ensuring high performance.
2.2 NIC Environments
We consider the scenario involving a group of GPU devices engaged in collaborative LLM training. These GPUs are from clusters equipped with different types of RDMA NIC, including InfiniBand or RoCE. We summarize the following two cases within this context.
  • Case 1: Homogeneous Clusters with High-Speed Interconnects. These GPU devices are distributed in clusters that have the same type of RDMA NIC, each equipped with high-speed interconnect (network switch). All GPU devices can communicate through RDMA NICs.
  • Case 2: Heterogeneous Clusters with Absence of High-Speed Interconnects. These GPU devices are distributed across multiple clusters, each equipped with either the same or distinct RDMA NICs, where high-speed interconnects are absent. Communication via RDMA NICs is confined to GPU devices within the same cluster, while GPU devices in different clusters communicate using Ethernet NICs.
2.3 Problem Description
We describe the core problem addressed in this paper as a scheduling problem for computation tasklets in the heterogeneous NIC environment. The distributed LLM training task is implemented using data and model parallelism techniques, where the entire computational task is decomposed into many parallel groups. Each group represents a set of parallel computations, with each computation viewed as a tasklet. Each tasklet corresponds to the forward and backward computations in a training iteration and is executed on certain GPU devices.
In the heterogeneous NIC environment, the training procedure is constrained by communication. The objective of Holmes is to achieve acceleration by allocating tasklets that require high communication volumes to GPU devices with faster connections. This is achieved by leveraging the characteristics of the netowrk environment, ensuring the maximization of device throughput in the LLM training process. The described method incorporates two parallelism techniques: pipeline and data. Pipeline parallelism involves the concurrent processing of computations in multiple stages. Each device is responsible for activation or gradient computation pertaining to different micro-batches in parallel, and the results can subsequently be communicated or transferred to subsequent stages. On the other hand, data parallelism implies that devices independently compute gradients for various micro-batches, necessitating synchronization of these gradients through communication.
We use the following two metrics.
  • TFLOPS stands for the achieved teraFLOP/s per GPU, representing the number of floating-point operations a GPU can perform in one second. It measures the GPU utilization rate, and the computational formula aligns with that in [25]. Further details are provided in the Experiment section (see Equation 6).
  • Throughout stands for the number of samples processed per second during the LLM training procedure. It measures the end-to-end training speed.
Figure 2: An illustration of parallelism mechanism in Holmes. In this case, a transformer-based model with 6 layers is trained across 2 clusters. Each cluster comprises 2 nodes, each equipped with 4 GPU devices. Communication between Node 1 and Node 2 utilizes InfiniBand, while Node 3 and Node 4 use RoCE. However, there is no high-speed interconnect between the two clusters, and communication between them relies solely on low-speed Ethernet. For parallelism settings, the degrees of data, tensor, and pipeline parallelisms are 2, 2, and 4, respectively. Pipeline parallelism is implemented between the two clusters using Ethernet. The model’s layers are unevenly partitioned into 2 stages and further distributed to different GPU devices across the two clusters. Data parallelism is performed within each cluster using RDMA, and tensor parallelism is implemented within each node using PCI-E.
2.4 Formalization
The scheduling problem is formalized as follows.
  • Let’s take into account a collection of $M$ clusters, which we denote as $\mathbf{C}=\{c_{1},c_{2},\ldots,c_{i},\ldots,c_{M}\}$. It is assumed that each cluster $c_{i}$ comprises $f_{i}$ nodes and that each node is equipped with $G$ devices. The total number of GPU devices is denoted as $N=G\cdot\sum_{i=1}^{M}f_{i}$.
  • We sequentially number clusters, nodes, and GPU devices. In the $i$-th cluster ($0<i\leq M$), the $j$-th GPU device ($0<j\leq G$) in the $k$-th node ($0<k\leq f_{i}$) is denoted as $rank_{G\cdot((\sum_{a=1}^{i-1}f_{a})+k-1)+j}$ in the global context.
  • The pipeline parallel degree is represented by $p\ (0<p\leq N/G)$, the tensor parallel degree by $t\ (0<t\leq G)$, and the data parallel degree by $d\ (0<d\leq N)$. It is important to note that the product of these degrees, $d\cdot p\cdot t$, is equal to $N$, which represents the total number of GPU devices.
  • For pipeline parallelism with a degree of $p$, there are a total of $t\cdot d$ pipeline parallel groups. Each of these groups can be represented as a matrix $[\textbf{PP}]\in\mathbb{Z}_{+}^{(t\ \cdot\ d)\times p}$. Similarly, for tensor and data parallelism with degrees $t$ and $p$, there are a total of $p\cdot d$ tensor parallel groups and $d\cdot t$ data parallel groups, respectively. These two types of group can be represented as matrices $[\textbf{TP}]\in\mathbb{Z}_{+}^{(p\ \cdot\ d)\times t}$ and $[\textbf{DP}]\in\mathbb{Z}_{+}^{(p\ \cdot\ t)\times d}$.
Different scheduling methods may lead to different parallel groups, each with its different communication cost. Our goal is to identify the most efficient scheduling method that minimizes the communication costs of the overall training procedure in the heterogeneous NIC environment.
3 Holmes Design
In this section, we present the Holmes design. Holmes is developed within the framework of Megatron-LM [32, 25], a popular tool for LLM training. Our approach re-optimizes two types of parallelism techniques: model parallelism and data parallelism. Model parallelism is further categorized into tensor parallelism and pipeline parallelism.
In distributed training, pipeline parallelism typically has less communication overhead than data parallelism because different GPU devices handle sequential computational stages with less inter-device communication between them [15, 23]. For data parallelism, data is divided into segments, and each device independently processes a portion. These units simultaneously perform identical operations on separate data, leading to increased communication for synchronization and greater overhead [28, 30]. Holmes enhances the built-in communication methods for pipeline and data parallelism in Megatron-LM. Figure 2 illustrates the parallelism mechanism in Holmes. Further details will be discussed in the subsequent parts.
3.1 Model Parallelism
3.1.1 Tensor Model Parallelism
Tensor model parallelism involves distributing individual layers of the model across multiple GPU devices. In the tensor parallelism of our framework, different layers of the model are assigned to various GPU devices within the same node, enabling concurrent processing of these layers, and thereby accelerating the training process. As a result, the communication operations associated with tensor parallelism predominantly take place within the same node, leveraging technologies such as PCI-E [16, 18] or NVLink [10] to facilitate inter-device communication.
Holmes adopts the same partitioning strategy used by Megatron-LM for transformer layers, which serve as the fundamental components of LLMs [25]. Following the partitioning process, we obtain the tensor parallel matrix denoted as $[\textbf{TP}]\in\mathbb{Z}_{+}^{(p\ \cdot\ d)\times t}$. This matrix contains various parallel groups, and the GPU devices within each group belong to the same node. We represent $[\textbf{TP}]_{i,j}$ as the $j$-th device in the $i$-th tensor parallel group, which can be formalized as follows:
${[\textbf{TP}]_{i,j}=rank_{(i-1)\cdot t+j}}\quad_{j\in\{1,2\ ,...,\ t\}}^{i\in\{1,2\ ,...,\ p\ \cdot\ d\}}$ (1)
Due to space limitations, we omit the further technical introduction in this section. For more detailed information on tensor parallelism, refer to [25].
3.1.2 Pipeline Model Parallelism
In traditional pipeline parallelism, the model layers are segmented into different computational stages, each assigned to different GPU devices. These devices execute the stages sequentially, transferring their results to subsequent stages. Communication operations occur through the exchange of data between these devices. When dealing with models that contain repeated transformer blocks, each GPU device can effectively manage an equal number of these blocks. This process involves breaking a batch into smaller micro-batches and executing them in a pipeline fashion [15, 23, 24]. We use periodic pipeline flushes to maintain the synchronization of optimizer steps between GPU devices.
The implementation of our pipeline parallelism is similar to PipeDream-Flush [24]. To support LLM training across clusters equipped with heterogeneous NICs, we make the following innovative improvements.
Cross-Cluster Pipeline Parallelism.
Current LLM training frameworks exclusively support LLM training with low-speed Ethernet NICs within clusters that possess heterogeneous NICs. To address these limitations, we have introduced a solution to enable LLM training in the heterogeneous NIC environment. Capitalizing on the relatively low communication cost of pipeline parallelism, we establish pipeline parallelism groups among multiple clusters that lack high-speed interconnects. Consequently, these pipeline parallelism groups consist of GPU devices connected via heterogeneous NICs, and their sole means of communication is through Ethernet due to NIC incompatibility. Conversely, owing to the division of pipeline parallelism groups, the data parallelism groups can comprise GPU devices connected via homogeneous NICs, enabling communication via RDMA NICs. In terms of tensor parallelism, it is confined to the same node and communicates via NVLink interfaces.
NIC Env TFLOPS Throughtout Bandwidth (Gbps)
InfiniBand 197 99.23 200
RoCE 160 80.54 200
Ethernet 122 61.32 25
Table 1: Comparison of TFLOPS and Throughput metrics when training a GPT model with 3.6 billion parameters on 4 nodes using InfiniBand, RoCE, and Ethernet respectively. Each node is equipped with 8 NVIDIA A100 GPUs. We also list the bandwidth of different NICs in the 3rd column.
Considering that data parallelism relies more on high-speed RDMA NICs than pipeline parallelism does, our Cross-Cluster Pipeline Parallelism method significantly enhances LLM training performance by scheduling pipeline parallelism groups across different clusters in heterogeneous NIC environments.
Self-Adapting Pipeline Partition.
Traditional pipeline parallelism was initially designed for training LLM in the homogeneous NIC environment. In such settings, the model layers are uniformly divided into different stages, ensuring relatively equal computational speeds across various stages of pipeline parallelism. This uniform partitioning is considered the most suitable choice, as all GPU devices exhibit equal computational speeds [25]. However, in the context of a heterogeneous NIC environment, different GPU devices may demonstrate varying computational speeds due to the diverse NICs to which they are connected. Table 1 presents a comparison of GPU performance using three different types of NIC during LLM training. Even if InfiniBand and RoCE NICs have the same bandwidth, the GPU device equipped with different types of NIC may exhibit significant variations in actual computational speed (TFLOPS). The observed disparities in TFLOPS and Throughput metrics underscore the inadequacy of a straightforward uniform partition for pipeline parallelism in the heterogeneous NIC environment.
In response to this challenge, we have developed a self-adaptive pipeline partitioning strategy that takes into consideration the specific NIC connected to each GPU device. This strategy optimizes the training process in the heterogeneous NIC environment by distributing varying numbers of layers in the model to different devices based on their connected NIC. To formalize the Self-Adapting Pipeline Partition strategy, we define the computational speed of a device within InfiniBand and RoCE as $S(\textsc{IB})$ and $S(\textsc{RoCE})$, respectively, interpreted as TFLOPS. Assuming a pipeline parallelism group with two devices connected to different NICs and a total number of layers for the model denoted as $N$, the number of layers distributed to the device connected to the InfiniBand NIC (denoted as $N_{ib}$) is calculated as follows:
$N_{ib}=\lfloor\frac{\alpha\cdot S(\textsc{IB})}{S(\textsc{IB})+S(\textsc{RoCE})}\cdot N\rfloor$ (2)
where $\alpha$ is the hyper-parameter. Similarly, the calculation of $N_{roce}$ follows the same method, ensuring $N_{ibz}+N_{roce}=N$. It can be easily extended to scenarios where the pipeline parallel size is more than 2. Our objective is to allocate a greater number of model layers to the GPU device connected to the faster NIC.
Summary.
We introduce two innovative enhancements to pipeline parallelism: optimizing communication operations and adapting to the heterogeneous NIC environment. We illustrate the process with a case presented in Figure 2.
To formalize the process of pipeline parallelism, we represent $[\textbf{PP}]_{i,j}$ as the $j$-th device in the $i$-th pipeline parallel group, which can be formalized as follows:
${[\textbf{PP}]_{i,j}=rank_{i+(j-1)\cdot t\cdot d}}\quad_{j\in\{1,2\ ,...,\ p\}}^{i\in\{1,2\ ,...,\ t\ \cdot\ d\}}$ (3)
The formulation is applicable to both homogeneous and heterogeneous NIC environments. In the heterogeneous environment, $[\textbf{PP}]_{i,*}$ represents the $i$-th pipeline parallel group comprising $p$ GPU devices that are connected to different types of NIC.
3.2 Data Parallelism
In traditional data parallelism, each GPU device possesses a complete replica of the model, while the input dataset is partitioned among these GPU devices. Periodically, these GPU devices synchronize their gradients to maintain consistent model weights across all of them. In cases where the model is too large to fit within a single worker, data parallelism is often used in conjunction with tensor and pipeline parallelism. In this scenario, data parallelism is applied to the smaller model segments that result from the process of model parallelism. The primary advantage of data parallelism lies in its ability to scale deep learning training, allowing the processing of larger datasets in less time. This scalability is particularly beneficial for training LLMs with large datasets.
Communication operations in data parallelism occur in the following steps:
  • Data Partition: The input dataset is divided into smaller batches, with each batch assigned to a different GPU for processing.
  • Gradient Computationn: Each GPU computes gradient information to help update the model parameters.
  • Gradient Aggregation: At the end of each batch or epoch, GPUs share and combine their calculated gradient information using the Ring Allreduce algorithm [31]. This algorithm circulates gradient information from each GPU to others in a circular manner, accumulating gradients on each GPU. The process is repeated through multiple steps until all GPUs have received the global gradient information.
  • Model Update: After receiving the combined gradient information, all GPUs use it to update the model parameters, ensuring that all GPUs have consistent model versions.
Communication between GPUs primarily takes place during the Gradient Aggregation step. This communication is facilitated by high-speed RDMA NIC and libraries like NCCL [4], ensuring the synchronization of gradients between GPUs and maintaining consistent model parameters across all devices.
Similarly to pipeline parallelism, traditional data parallelism also does not take into account the heterogeneous NIC environment, and it can only support using the low-speed Ethernet NIC when training LLMs in the heterogeneous environment. To address the problem and accelerate the data parallelism in both homogeneous and heterogeneous environments, we make the following innovative improvements.
Automatic NIC Selection.
We have implemented an auto-select strategy by modifying NCCL [4] and Megatron-LM [3] to support hybrid setups of NICs, including InfiniBand, RoCE, and Ethernet. Our strategy provides functions that create data parallelism groups based on the actual network environment, ensuring that GPU devices within the same data parallelism group are connected to the same type of NIC.
Traditional data parallelism techniques do not ensure the allocation of all data parallelism groups to GPU devices sharing the same type of NIC. In scenarios where a data parallelism group includes two GPU devices linked to InfiniBand and RoCE NICs, communication between the two devices is limited to Ethernet due to their inherent incompatibility. Consequently, traditional data parallelism frequently faces performance bottlenecks, particularly during the Gradient Aggregation step, as it necessitates waiting for all GPU devices to conclude their computations. Detailed comparative results are presented in the subsequent Experiment section.
Overlapped Distributed Optimizer.
Building upon recent research [2], we have incorporated a novel partition strategy for gradients and optimizer states into our LLM training framework. This strategy, known as the OverlappedDistributedOptimizer, can be summarized by the following core principles:
  • Traditional optimizers such as Adam and SGD independently update individual parameter values, making it unnecessary to maintain parameters as complete entities.
  • Each collective communication operation should transmit a significant amount of data to optimize the utilization of communication bandwidth.
  • There should be no additional communication overhead or GPU memory copying involved.
The OverlappedDistributedOptimizer presents an effective approach to seamlessly intertwining communication with computation. This strategy maximizes the potential for overlapping and utilizes interconnection bandwidth to its fullest extent, all while avoiding any additional overhead in terms of communication volume. For in-depth technical details, you can refer to [2]. Specifically, the OverlappedDistributedOptimizer is orthogonal to our previous optimizations and can be combined to further enhance training efficiency.
Parameter
Group
Number of
Parameters (billion)
Attention
Heads
Hidden
Size
Number
of Layers
Tensor
Parallel Size
Pipeline
Parallel Size
Mircro
BatchSize
Batch
Size
1 3.6 32 3072 30 1 2 4 768
2 1536
3 7.5 4096 36 1536
4 2688
5 3 1536
6 2688
7 39.1 64 8192 48 8 2 1536
8 3
Table 2: Details of parameter groups for training GPT models ranging from 3.6 billion to 39.1 billion parameters. Since our optimization focuses on data parallelism and pipeline parallelism, without involving tensor parallelism, for Parameter Group 1 to Parameter Group 6, we set the tensor parallel size to 1. For Parameter Group 7 and 8, due to the large parameter size of the model, we set the tensor parallel size to 8.
Summary.
We introduce two innovative enhancements to data parallelism: an optimal selection strategy of NICs in the heterogeneous NIC environment and a well-designed distributed optimizer. We illustrate the process with a case presented in Figure 2.
To formalize the process of data parallelism, we represent $[\textbf{DP}]_{i,j}$ as the $j$-th device in the $i$-th data parallel group, which can be formalized as follows:
${[\textbf{DP}]_{i,j}=rank_{mod(i-1,t)+(floor(\frac{i-1}{t})\cdot d+j-1)\cdot t+1}}\ _{j\in\{1,2\ ,...,\ d\}}^{i\in\{1,2\ ,...,\ p\ \cdot\ t\}}$ (4)
where $mod(i-1,t)$ denotes the modulo operation, which calculates the remainder when dividing two numbers, and $floor(*)$ is used to signify rounding down a real number.
4 Experiment
In this section, we seek to answer the following questions:
Q1: How well does Cross-Cluster Pipeline Parallelism perform for models of different sizes?
Q2: How well does Automatic NIC Selection perform in the context of data parallelism?
Q3: How well does Self-Adapting Pipeline Partition perform compared with Uniform Pipeline Partition?
Q4: How well does Holmes perform compared to other mainstream LLM training frameworks?
Q5: How well do the different components of Holmes contribute to LLM training?
Parameter Group NIC Env 4 Nodes / 32 GPUs 6 Nodes / 48 GPUs 8 Nodes / 64 GPUs
TFLOPS Throughout TFLOPS Throughout TFLOPS Throughout
1 InfiniBand 197 99.23 188 142.09 148 148.88
RoCE 160 80.54 151 114.15 145 145.64
Ethernet 122 61.32 99 74.98 83 83.38
Hybird 149 74.91 129 97.84 112 112.46
2 InfiniBand 206 103.66 200 151.25 156 156.66
RoCE 168 84.78 162 122.53 159 160.47
Ethernet 145 72.95 128 96.75 114 114.52
Hybird 162 81.38 152 114.63 132 132.73
3 InfiniBand 229 55.95 220 80.64 189 92.35
RoCE 196 48.04 185 67.84 185 90.40
Ethernet 168 41.04 143 52.91 132 64.85
Hybird 191 46.66 170 62.43 168 82.02
4 InfiniBand 233 57.03 228 83.61 196 95.79
RoCE 201 49.10 193 70.88 194 94.85
Ethernet 180 44.10 168 61.59 158 77.31
Hybird 200 48.89 187 68.52 177 86.58
Table 3: Performance of different parameter groups in both homogeneous and heterogeneous NIC environments. Here the Hybird environment contains two clusters with the same number of nodes.
Figure 3: The time cost of grads-reduce-scatter operation with different parameter groups in both homogeneous and heterogeneous NIC environments.
Parameter Group NIC Env 6 Nodes / 48 GPUs 12 Nodes / 96 GPUs
2RoCE & 2RoCE & 2IB 2RoCE & 2IB & 2IB 4RoCE & 4IB & 4IB
TFLOPS / Throughout TFLOPS / Throughout
5 Ethernet 143 / 52.51 95 / 70.11
Hybird 163 / 59.75 161 / 59.19 138 / 101.24
6 Ethernet 160 / 59.59 122 / 89.65
Hybird 174 / 63.96 169 / 61.87 146 / 107.21
Table 4: Performance of different parameter groups in the heterogeneous NIC environment. Here the Hybird environment contains three clusters with the same number of nodes.
Figure 4: Throughput of different parameter groups in training GPT model on 4 nodes within both homogeneous and heterogeneous NIC environments.
4.1 Experiment Setup
The experimental setup encompasses the following aspects.
Machine Environment. All of our results are run with mixed precision. Each cluster node is equipped with 8 NVIDIA A100 GPUs, each with 80 GB of memory [5], interconnected by NVLink [10]. The A100 GPUs achieve a peak device throughput of 312 teraFLOP/s with 16-bit precision. For the comparison of our results, we report results including the achieved teraFLOP/s per GPU (TFLOPS) and aggregate Throughput, calculated by multiplying the throughput of each individual GPU.
NIC Environment. To simulate different cases discussed in Section 2.2, we conducted experiments across different types of NIC environments. The characteristics of these NIC environments are summarized below.
  • InfiniBand: The cluster nodes are equipped with high-speed InfiniBand NICs.
  • RoCE: The cluster nodes are equipped with high-speed RoCE NICs.
  • Ethernet: The cluster nodes lack high-speed RDMA NICs and can only communicate through Ethernet NICs.
  • Hybird: There are multiple different clusters, each with nodes equipped with different high-speed NICs.
Model Selection. For our experiments, we employ GPT models of different sizes. Specifically, the model must be compatible with the number of model-parallel GPUs used in the experiment. We utilize standard model architectures, such as GPT-3 [6] when appropriate.
Parameter Group. To demonstrate the end-to-end performance of Holmes in training GPT models of various sizes in the heterogeneous NIC environment, we defined several parameter groups. Table 2 provides detailed information on the parameter groups used in our experiments. Specifically, we employ the interleaved pipeline schedule with scatter/gather optimization enabled. All models utilize a vocabulary size ($V$) of 51,200 (a multiple of 1024) and a sequence length ($s$) of 2048. We vary the hidden size ($h$), the number of attention heads, and the number of layers ($l$). The number of parameters in a model, denoted by $P$, can be computed as follows:
$p=12lh^{2}(1+\frac{13}{12h}+\frac{V+s}{12lh})$ (5)
As the model size increases, we concurrently augment the batch size ($B$) and the number of GPUs ($n$). The predominant portion of floating-point operations within the model occurs in the matrix multiplications (GEMMs) within the transformer and logit layers. When focusing solely on these GEMMs, the number of FLOPs per iteration is calculated as follows:
$F=96Bslh^{2}(1+\frac{s}{6h}+\frac{V}{16lh})$ (6)
This serves as a lower bound for the actual FLOP count but is expected to be closely aligned with the true value. We consider a FLOP as a floating-point operation, irrespective of precision [25].
In subsequent parts, we evaluate each parameter group with varying numbers of nodes under different NIC environments.
Hyper-Parameter and Compared Frameworks. The experimental results of Holmes presented in Figure 5, 6, 7, and Table 5 are based on the Self-Adapting Pipeline Partition strategy, with the hyperparameter $\alpha$ set to 1.05. The LLM training frameworks compared with Holmes include Megatron-LM [3], Megatron-LLaMA [2], and Megatron-DeepSpeed [1].
4.2 Performance Comparison
Different NIC Environments.
As discussed in 2.2, we compare two cases in the heterogeneous NIC environment. For Case 1 and Case 2, we evaluate Holmes across four types of NIC environments, including both homogeneous and heterogeneous high-speed interconnects, and varying numbers of nodes. The results are presented in Table 3, providing insights to address questions Q1 and Q2.
By analyzing the results in Table 3, it becomes evident that, with the integration of two components, Cross-Cluster Pipeline Parallelism, and Automatic NIC Selection, Holmes demonstrates the ability to efficiently train LLM in Hybrid environment. It achieves performance levels that closely resemble those attainable in the homogeneous InfiniBand and RoCE environments, and significantly surpasses training efficiency in Ethernet environment.
We also present the time cost of a specific operation during LLM training in various NIC environments. Figure 3 provides a comparative result of the time cost associated with the reduce-scatter operation, a typically time-consuming process in data parallelism. As the choice of NIC can significantly influence the time cost of this operation, the results in Figure 3 demonstrate that within the homogeneous InfiniBand environment, reduce-scatter exhibits the shortest duration, aligning with the results in Table 3 where LLM training achieves the highest TFLOPS and Throughput in the homogeneous InfiniBand environment. Furthermore, the results in Hybrid and Ethernet environments indicate that Holmes optimally utilizes the heterogeneous RDMA NICs, effectively reducing the time cost of reduce-scatter and achieving faster training speed.
Specially, for Case 2, we evaluated Holmes in a scenario where high-speed interconnects are entirely absent between GPU clusters. This situation is common when multiple GPU clusters are distributed across different locations. Figure 4 presents the results, demonstrating Holmes’s performance in a scenario involving LLM training across multiple GPU clusters without high-speed interconnects (denoted as InfiniBand & Ethernet and RoCE & Ethernet). In this context, Holmes employs pipeline parallelism across two clusters with Ethernet NICs, allowing data parallelism to fully leverage high-speed RDMA NICs. To provide a clear efficiency comparison of Holmes, we present performance results for GPU clusters with high-speed interconnects (denoted as InfiniBand and RoCE). This serves as an upper bound for throughput. Additionally, we present performance results for GPU clusters with only slow-speed interconnects (denoted as Ethernet), serving as a lower bound for throughput. We compare the performance results of Holmes in clusters with heterogeneous NICs on the same scale (denoted as Hybrid) for reference. The results in Figure 4 demonstrate that Holmes achieves competitive training performance in a scenario involving multiple GPU clusters without high-speed interconnects, regardless of the heterogeneity or homogeneity of the NICs between clusters.
We conducted additional evaluations of Holmes in a more intricate scenario, involving three GPU clusters with an equal number of nodes but varying NIC types. The results are detailed in Table 4. In this context, Holmes utilizes pipeline parallelism with a degree of 3, effectively partitioning the model into three stages. The results demonstrate that Holmes consistently maintains high performance, even when dealing with a context that includes more than two clusters. This highlights the adaptability of Holmes for training LLM across multiple GPU clusters.
Figure 5: Performance of different parameter groups in training GPT model with various pipeline partition strategies.
Pipeline Partition.
To answer question Q3, we conducted a comparison between Self-Adapting Pipeline Partition and Uniform Pipeline Partition. The results are presented in Figure 5. The results substantiate our hypothesis that Uniform Pipeline Partition is not the optimal choice for training LLM in the heterogeneous NIC environment. Holmes employs Self-Adapting Pipeline Partition to account for variations in training speeds between two GPU clusters, resulting in higher TFLOPS and Throughput during the LLM training process compared to traditional Uniform Pipeline Partition strategies.
Comparison with Other Training Frameworks.
To answer question Q4, we conducted a comparative analysis on the performance of Holmes against other prominent LLM training frameworks, namely Megatron-LM, Megatron-DeepSpeed, and Megatron-LLaMA, within the heterogeneous NIC environment. The results of this comparison are presented in Figure 6. Notably, Holmes outperforms the other LLM training frameworks, emerging as the top performer. This superiority can be attributed to Holmes being the sole training framework specifically designed for heterogeneous NIC environments. Significantly, Megatron-LLaMA demonstrates superior performance compared to Megatron-LM and Megatron-DeepSpeed, highlighting the effectiveness of Overlapped Distributed Optimizer component proposed by the former.
Figure 6: Performance comparison between Holmes and other mainstream LLM training frameworks: experiment conducted using parameter group 3 on 8 nodes (4 nodes equipped with RoCE NICs and 4 nodes equipped with IB NICs)
Furthermore, we also evaluate the speedup ratios of Holmes in comparison to these frameworks. The result is displayed in Figure 7. As expected, Holmes achieves the highest speedup compared to other LLM training frameworks, which further demonstrates the efficiency in training scalability. We believe that Holmes can scale to distributed training involving tens to hundreds of nodes.
4.3 Ablation Study
As illustrated in Figure 1, the Holmes Framework comprises key components designed to enhance the training speed of LLMs in the heterogeneous NIC environment. These components include Cross-Cluster Pipeline Parallelism, Self-Adapting Pipeline Partition, Automatic NIC Selection, and Overlapped Distributed Optimizer.
To answer question Q5 and investigate the impact of different components in Holmes, we conducted an ablation study in this subsection. The effect of Cross-Cluster Pipeline Parallelism has been validated in the previous subsection, and here we focus on the remaining three components. The results of the ablation study are shown in Table 5. From the results, we observe that the component of Overlapped Distributed Optimizer contributes more than the component of Self-Adapting Pipeline Partition. Furthermore, from the results in the last column, it is evident that the effects of Overlapped Distributed Optimizer and Self-Adapting Pipeline Partition are nearly orthogonal. This suggests that these two components enhance the training speed from different aspects.
By removing the components of Overlapped Distributed Optimizer and Self-Adapting Pipeline Partition from Holmes, we can evaluate the impact of Automatic NIC Selection. The results in the first and last lines of Table 5 demonstrate that even with only the Automatic NIC Selection component, the effectiveness of Holmes surpasses that of the Megatron-LM framework by a significant margin.
Figure 7: Comparison of speedup ratios in training GPT model with varying numbers of nodes: Holmes versus other mainstream training frameworks. The experiment was conducted using Parameter Groups 7 and 8.
Training Framework TFLOPS Throughout
Megatron-LM 132 64.86
Holmes 183 89.48
w/o Self-Adapting-Partition 179 (-4) 87.55 (-2.43)
w/o Overlapped Optimizer 170 (-13) 83.15 (-6.83)
w/o Above Two 168 (-15) 82.02 (-7.96)
Table 5: Ablation study of different components in Holmes. The experiment is conducted using the same setting as Figure 6.
4.4 Discussion
In this section, we undertake comprehensive experiments from several aspects to address the questions posed at the beginning of this section. Firstly, we evaluate the performance of Holmes in various NIC environments, constituting a pivotal experiment within this paper. Our findings demonstrate that Holmes attains a performance level that competes favorably in heterogeneous NIC environments, closely approaching the efficiency achievable in homogeneous RDMA-capable networks (InfiniBand or RoCE), and surpassing the training efficiency observed in purely Ethernet environments. Subsequently, we conduct an in-depth comparative analysis of the different components of Holmes, substantiating their effectiveness. Finally, we compare Holmes with other mainstream LLM training frameworks, thereby verifying its advancements.
5 Related Work
5.1 Large Language Models
Large language models (LLMs) have attracted considerable attention in recent years, marking a substantial advancement in the field of NLP. Their origins can be traced back to the introduction of the Transformer architecture [36], which, relying on self-attention mechanisms, brought about a revolution in deep learning, particularly in the realm of understanding and generation of natural languages. Notable examples of LLMs include the GPT series [22] and BERT [8]. The fundamental concept underlying LLMs is pre-training on extensive text corpora, which equips the models with a broad grasp of language, encompassing syntax, semantics, and world knowledge. Subsequently, these models undergo fine-tuning for specific tasks, leading to state-of-the-art performance in a series of NLP tasks like language modeling, text classification, machine translation, and question-answering.
5.2 Distributed Training Frameworks
Distributed training frameworks are essential for efficient and scalable Large Language Model (LLM) training. They enable parallel task execution across multiple devices, addressing computational and memory demands of large models. LLMs are usually trained in data centers with homogeneous high-speed RDMA NICs connecting GPUs. However, existing frameworks often neglect training LLMs in heterogeneous NIC environments.
Some popular distributed training frameworks tailored for Large Language Models (LLMs) include Megatron-LM [32, 25, 3], Megatron-LLaMA [2], and DeepSpeed [27, 30, 28, 1]. Megatron-LM, developed by NVIDIA, efficiently handles training of very large models using a combination of model parallelism and data parallelism across multiple GPUs and nodes. This approach enables LLM training by sharing model parameters among GPUs and processing data in parallel [25, 39]. Megatron-LM excels in minimizing synchronization delays during training through optimized communication between nodes and GPUs, utilizing efficient communication and synchronization techniques. Megatron-LLaMA, an extension of Megatron-LM, enhances the framework by integrating a standardized LLaMA implementation and introducing an efficient communication-computation parallelism method. Another notable framework is DeepSpeed, an open-source library from Microsoft designed for efficient training of large-scale deep learning models. DeepSpeed features memory-efficient training, accommodating models larger than GPU memory capacity through techniques like mixed-precision training, gradient accumulation, and offloading optimizer states. Its ZeRO optimization technology further improves memory efficiency by distributing model weights and optimizer states among GPUs and nodes, reducing memory usage and enabling training of even larger models [27, 30].
5.3 Communication Optimization
Distributed training necessitates frequent parameter synchronization across nodes. In the initial stages of communication scheduling, the primary focus is on ensuring model consistency [35, 29, 14] and refining parameter synchronization architectures [19, 26, 12]. With the increasing scale of models, certain research endeavors explore coordinated optimization of both communication and parallel strategies [37, 43]. This involves implementing advanced strategies to enhance parallel efficiency while concurrently reducing communication demands.
Recent advancements in distributed training efficiency, exemplified by TopoOpt [37], focus on coordinating improvements in computation, communication, and network topology. TopoOpt employs alternative optimization techniques and draws inspiration from group theory, utilizing the TotientPerms algorithm to optimize network topology, routing plans, and parallelization strategies. Another noteworthy project, Alpa [43], addresses challenges in cross-mesh sharding by introducing a broadcast-based communication mechanism and an overlap-friendly pipeline scheduling strategy.
6 Conclusion
In this paper, we introduce Holmes, a meticulously designed framework designed for training LLM across multiple GPU clusters. Our empirical studies indicate that the proposed training framework can achieve performance levels in the heterogeneous NIC environment comparable to those possible with homogeneous RDMA NICs. Holmes is considered a significant advancement in making LLM training more accessible, thereby expanding the potential for efficient scaling within the broader research community.
7 Acknowledgement
This work is supported by National Natural Science Foundation of China Grant (No. U22A6001) and Key Research Project of Zhejiang Lab (No. 2022PG0AC02).